# Foundational Policy Acquisition via Multitask Learning for Motor Skill Generation

### v3

@@ IV Experiments
We evaluated the proposed learning method using the monopod heading task, as shown in Fig. 1. Simulated experiments were conducted on the sagittal plane using a dynamics simulator MuJoCo [28]. Through this dynamic movement control task, we showed that the monopod robot model can adaptively accomplish multiple tasks in multiple environments using the proposed method.
IV-A Monopod heading task
In the monopod heading task, the robot model tried to hit a ball back to a specified goal point while maintaining its balance until it hit the ball. As small changes in the movement result in significantly different reward values, precise dynamic movement control is necessary. The reward function for the task is defined as follows:
@@ IV-A Monopod heading task
where $p_{\mathrm{ball}}$ denotes the ball position; $p_{\mathrm{targ}}(t,c)$ is the context-dependent desired ball trajectory; and $\tau$ denotes the joint torque vector. The desired ball trajectory was derived using the motion equation from a desired hit position to a target point within a given ball movement duration. The learning trial was terminated if the ball position deviated from the desired ball trajectory by a certain distance. The reward parameters were set to $\lambda_{1}=10.0$ and $\lambda_{2}=5.0\times 10^{-3}$. Furthermore, we terminated the learning trial when the distance between the head position and the desired hitting position exceeded a certain threshold, primarily because the monopod robot fell to the ground.
The monopod robot model had five joints, each controlled by a proportional servo (P) controller. The state variables of the robot are listed in Table I. Notably, the robot did not observe the goal position, although the task achievement was informed through the reward signal. The action variables were the target joint angles for the P controller at each joint.
TABLE I: Observables
@@
In this task, five context variables were considered, which are presented in Table II. The coefficient of restitution defines a physical parameter of the ball and describes the relationship between ball velocities before and after hitting it. Horizontal and vertical goal position variables define different goal locations, and horizontal and vertical throw-in positions represent different ball throw positions toward the monopod robot.
TABLE II: Context variables
@@
IV-B Network models
The policy network was implemented using the multilayer neural network model. The network model had two hidden layers, shown in Fig. 2. The input variables to the network included the robot state $s$ and the latent variable of the encoder $z$. The encoder network encoded the context variable $c$ to the latent variable $z$ and implemented a neural network model with two hidden layers. In the policy search phase, the encoder and Q-function networks were trained using actual context variables $c$. Subsequently, in the adaptation phase, the encoder network was not used. The latent variable $z$ was estimated through adaptation trials.
IV-C Foundational policy model
In the foundational policy acquisition phase, we trained the networks under the combination of context variables. In other words, we conducted domain randomization with 32 different combinations of context variables as five context variables with two settings were considered, that is, $322^{5}=$ 32. We repeated this learning process 25 times to generate 25 different domain randomized policies with different random seeds.
IV-D Selecting the foundation policy network
@@
IV-E Evaluating the skill generation performance
We evaluated how the monopod model can efficiently generated skills to cope with novel tasks and environments using our proposed method. The parameters representing the novel settings were uniformly sampled from intervals defined in Table II. Concretely, we randomly sampled $32$ settings and fixed these values to evaluate the task performances of each method. We confirmed that these $32$ settings differed from those used in the foundational policy acquisition phase described in Section IV-C. We then compared the proposed approach with the policy acquired using the domain randomization method, which is widely used to cope with an unknown environment.
We then further explored the potential of our proposed approach. We set the goal position in the rear region of the robot, while all training in the foundational policy acquisition phase was done with the goal position in the front region. Thus, here we tried to evaluate the extrapolation performance of our skill generation method.
IV-F Evaluating policy selection performance
We compared the proposed policy selection method with random selection in terms of top-one regret, which has been adopted to evaluate policy models [29]. The top-one regret evaluates the error between the best return value and the actual return. In this study, we compared the return value of the selected policy using the proposed method with that of the best policy that acquired the highest return among the candidate policies as follows:
@@ IV-F Evaluating policy selection performance
where $R_{\mathrm{best}}$ denotes the return value of the best policy, and $R_{\mathrm{selected}}$ denotes that of the selected policy using the proposed method. To reliably evaluate the top-one regret, we used a bootstrap method. Concretely, as described in Section IV-C, we generated $25$ candidate policies. Subsequently, we sampled five policies with replacements from the $25$ candidates. Through the sampling procedure described above, we made 1000 sets of the five policies, that is, $N=1000$.
@@ V Results
V-A Skill Generation Performance
Domain randomization is a frequently used approach to cope with multiple tasks and environments. We first evaluated the acquired monopod heading policies using domain randomization as the baseline. We trained the networks under the combination of context variables. The procedure to train the network was the same as that described in Section IV-C. Notably, the proposed method also trained the encoder and policy networks, while the domain randomization method did not.
Fig. 3 and Fig. 3 show the monopod heading task performances at a novel goal position. As shown in the rightmost panel of Fig. 3, the ball deviated from the goal position when a policy acquired by domain randomization was adopted. Using the proposed method, the ball hit by the monopod successfully tracked the desired trajectory and passed around the goal position, as depicted in Fig. 3. Notably, the monopod model also learned to maintain balance when hitting the ball.
Fig. 3 shows the task performances of the policies trained by the domain randomization (green bar) and the proposed adaptation (red bar). Each box plot was depicted based on $32$ task performances, as described in Section IV-E. Both the domain randomization and generation phase of the proposed method trained the networks for $3\times 10^{6}$ time steps, which approximately corresponds to $45,000$ trials. The skill-generation phase only required additional $k_{\mathrm{max}}=100$ trials to perform novel tasks successfully.
@@ V-A Skill Generation Performance
Figure 3: Monopod heading performances with the novel goal position. The leftmost snapshot shows the timing of hitting the ball, that is, $t=1.0$ s. The green sphere indicates the goal position. The red dot indicates the desired trajectory. (a) Heading performance by domain randomization. As shown in the rightmost panel, at $t=2.0$ s, the ball deviated from the goal position. (b) Heading performance is performed using the proposed method. The hit ball successfully tracked the desired trajectory and passed around the goal position. Notably, the monopod also learned to maintain its balance with the hit ball. (c) Skill generation performance: With Domain Randomization Multitask heading performance. The domain randomization approach could not achieve a satisfactory multitask performance (green bar). Multiple tasks in multiple environments are accomplished using the proposed method (red bar).
V-B Policy selection performance
Here, we evaluate how the policy model selected based on the proposed method performed better than a randomly selected model. We adopted the top-one regret as the performance measure to compare the selection performances. Section IV-F describes a concrete calculation of the top-one regret.
The top-one regret score of the proposed method was $1.4$, whereas that of the random selection was $5.1$. As the smaller regret showed better performance, we confirmed that the proposed method contributed to selecting a potentially well-performing policy after the adaptation phase.
V-C Novel skill generation performance
@@
We depicted how the proposed skill generation method improved the task performance in Fig. 4. As shown in Fig. 4, the hit ball did not track the desired trajectory and deviated from the goal before the adaptation. To evaluate the task performance before the skill generation, we set the latent variable to zero, that is, $z=0.0$, assuming that we did not have prior knowledge of the novel task before the skill generation. After the skill generation, the optimized skill successfully hit the ball towards the new goal position, as shown in Fig. 4.
We also evaluated the overall performances of the policies before and after the skill generation on $32$ different novel tasks, wherein the tasks were sampled randomly, as described in Section IV-E. Fig. 4 shows the task performances of each policy. Before the skill generation, for most novel tasks, task performances were unsatisfactory. Furthermore, performances were varied. After the skill generation, the generated policy for novel tasks consistently achieved successful heading performances.
Furthermore, when we placed the goal position at the back of the robot, the robot model generated a whole new motor skill to ’kick’ the ball backwards to the goal position instead of heading it. Figure 6 illustrated the generated overhead kick movement to hit the ball to the totally new goal position.
The above results indicate that the proposed low-dimensional skill representation and derivative-free optimisation successfully facilitate skill emergence in novel situations.
@@
Figure 4: Multitask Heading performance on novel tasks. The green sphere indicates the goal position. The red dot indicates the desired trajectory. (a) Foundational policy: the hit ball did not track the desired trajectory and deviated from the goal. (b) Generated skill: the adapted policy successfully hit the ball toward the new goal position. (c) Multitask heading performances: With the foundational policy, task performances were unsatisfactory and varied. With the generated skill, generated skills for novel tasks consistently achieved successful heading performances.
@@
(b) Skill generation perofrmance
Figure 5: (a)Fine-tuning performance. We found that task performances became worse than those before fine-tuning trials. This result indicated that the policy modal could not adapt to new domains with only a small number of additional trials. (b)The proposed method was compared with the baseline method using context variable information. Interestingly, the proposed skill-generation method showed a better heading performance than the method using context information. This result empirically showed that we extracted a meaningful latent space for the given movement category.
Figure 6: Generated overhead kicking skill. We set the goal position (green) in the rear region of the robot, while all training in the foundational policy acquisition phase was done with the goal position in the front region. Thus, here we tried to evaluate the extrapolation performance of our skill generation method. As the result, the robot model generated a whole new motor skill to ’kick’ the ball backwards to the goal position instead of heading it.
@@ VI-A Skill generation through latent variable optimization
Fig. 5 shows the heading performances with the $32$ randomly selected domains. We found that the task performances became worse than those before the fine-tuning trials. This result indicated that the domain randomization policy model could not adapt to new domains with only a small number of additional trials. In other words, extracting the low-dimensional latent space and optimizing the latent variable were significant procedures for generating a skill to cope with novel tasks and environments.
@@ VI-B Using explicit context information
In this study, we assumed that the policy could not observe the context variable in the adaptation phase as we usually lack information regarding novel domains. However, it would be interesting to investigate how task performance can be improved if the context variable was observable. Here, we explicitly encoded the context variable of $32$ novel domains using the trained encoder in the skill-generation phase. The encoder outputs were forwarded to the policy network as latent variables.
Fig. 5 shows the task performances using explicit context information. Interestingly, the proposed adaptation method showed better heading performances than the method using context information. On the one hand, this is because the context variables for novel domains were not used when the encoder network was trained. On the other hand, even the encoder used in our method was not trained with novel context information. However, our proposed method achieved better performance than the method that considers the context variables as inputs. This result empirically showed that we extracted a meaningful latent space for the given movement category.
@@ VII Conclusion
In this study, we proposed a foundational policy acquisition method via the multitask reinforcement learning approach. We particularly focused on dynamic movement tasks in which small movement changes lead to completely different consequences. We consider the multi-heading task to be concrete using the monopod model. The monopod learned to hit a ball back to multiple target positions in this task. Furthermore, the coefficient of restitution of the ball also varied. In the skill-generation phase of our proposed method, context variables such as goal positions and environmental settings were not provided to the monopod model. Therefore, the learning system needed to adapt to implicitly changed contexts. After the skill acquisition process, we showed that the monopod model accomplished $32$ different heading tasks. For the skill acquisition, we selected one foundational policy among $25$ candidate policies based on the return values acquired in the foundational-policy acquisition phase. These policy acquisition and selection strategies to find foundational policy models were important factors for successful skill generation. We further demonstrated that the robot model successfully generated a whole new motor skill to ’kick’ the ball backward to the goal position instead of heading it.
As we focused on a dynamic movement task that only requires a small change to accomplish different targets in different environmental settings, we could expect that the latent variable which represents different contexts can belong to a low-dimensional space. We empirically found that optimizing a one-dimensional latent variable to cope with different situations can effectively achieve higher return values.
As a future study, we aim to explore the possibility of the proposed method on a multi-agent learning problem. We are interested in observing how our proposed method can modify agent policy according to implicit changes in other agents’ strategies.

### v4

@@ IV Experiments
To evaluate the multitask learning performance, we first compared the proposed framework with previous meta-RL methods in the standard multi-locomotion domains [8, 2] using the Half-Cheetah model. We then applied the proposed method to the ball heading task using a monopod robot model to evaluate skill generation performance. We adopted a dynamics simulator MuJoCo [32] for these evaluations.
IV-A Multi-locomotion Domains
We compared our proposed approach with previous methods in the Half-Cheetah-Dir and Half-Cheetah-Vel domains. In the Half-Cheetah-Dir domain, the goal is to move forward or backward as fast as possible. In the Half-Cheetah-Vel domain, the goal is to reach a target velocity [8, 2]. In the policy acquisition phase, one of the target directions was selected as a context value in the Half-Cheetah-Dir domain, and the target velocity was sampled from a uniform distribution with a range of $0.0$ to $3.0$ m/s [2] as a context value in the Half-Cheetah-Vel domain. These context values are fixed for the duration of one learning episode, where the learning agents collected data for $200$ time steps in one episode. We trained the $K=5$ candidate policies and selected the best policy from these candidates by following the Algorithms 1 and 2. In the Half-Cheetah-Dir domain, the number of tasks was two, forward and backward. Therefore, the size of the context was $|\mathcal{C}|=2$. In the Half-Cheetah-Vel domain, one hundred tasks were randomly sampled from the uniform distribution, thus, the size of the context was $|\mathcal{C}|=100$. In the skill generation phase, we set the maximum number of iterations as $k_{\mathrm{max}}=100$. Table III in the Appendix lists the hyperparameters used in our proposed learning framework.
IV-B Monopod Heading Task
In the monopod heading task, the robot model tried to hit a ball back to a specified goal point while maintaining its balance until it hit the ball, as shown in Fig. 2(c) (d) (e). As small changes in the movement result in significantly different reward values, precise dynamic movement control is necessary. The reward function for the task is defined as follows:
@@ IV-B Monopod Heading Task
where $p_{\mathrm{ball}}$ denotes the ball position; $p_{\mathrm{targ}}(t,c)$ is the context-dependent desired ball trajectory; and $\tau$ denotes the joints torque. The desired ball trajectory was derived using the motion equation from a desired hit position to a target point within a given ball flying duration. The learning trial was terminated if the ball position deviated from the desired ball trajectory by a certain distance. The reward parameters were set to $\lambda_{1}=10$ and $\lambda_{2}=5\times 10^{-3}$. Furthermore, we terminated the learning trial when the distance between the head position and the desired hitting position exceeded a certain threshold, primarily because the monopod robot fell to the ground.
The monopod robot model had five joints, each controlled by a proportional (P) servo controller. The state variables of the robot are listed in Table I. Notably, the robot did not explicitly observe the goal position. Instead, the reward signal informed it of the task achievement. The action variables were the target joint angles for the P servo controller at each joint.
TABLE I: Observables of monopod robot model
@@
In this task, five context variables were considered, which are presented in Table II. The coefficient of restitution defines a physical parameter of the ball and describes the relationship between ball velocities before and after hitting it. Horizontal and vertical goal position variables define different goal locations, and horizontal and vertical throw-in positions represent different ball throw positions toward the monopod.
TABLE II: Context variables of monopod robot model
@@
IV-B1 Context randomization
In the foundational policy acquisition phase, we trained the networks under the combination of context variables. In other words, we conducted context randomization with 32 different combinations of context variables as five context variables with two settings were considered, that is, $322^{5}=$ 32. We repeated this learning process 25 times to generate 25 different domain randomized policies with different random seeds.
IV-B2 Selecting the foundation policy network
@@
IV-B3 Evaluating the skill generation performance
We evaluated how the monopod model can efficiently generate skills to cope with novel tasks and environments using our proposed method. The parameters representing the novel settings were uniformly sampled from intervals defined in Table II. Concretely, we randomly sampled the $32$ context and fixed these values to evaluate the task performances of each method. We confirmed that these $32$ settings differed from those used in the foundational policy acquisition phase described in Section IV-B1. We then compared the proposed approach with the policy acquired using the domain randomization method, which is widely used to cope with an unknown environment.
We then further explored the potential of our proposed approach. We set the goal position in the rear region of the robot, while all training in the foundational policy acquisition phase was done with the goal position in the front region. Thus, we tried to evaluate the extrapolation performance of our skill generation method here.
IV-B4 Evaluating policy selection performance
We compared the proposed policy selection method with random selection in terms of top-one regret, which has been adopted to evaluate policy [33]. The top-one regret evaluates the error between the selected return value of the chosen policy and the best return that acquired the highest return among the candidate policies as follows:
@@ IV-B4 Evaluating policy selection performance
where $R_{\mathrm{best}}$ denotes the return value of the best policy, and $R_{\mathrm{selected}}$ denotes that of the selected policy using the proposed method. To reliably evaluate the top-one regret, we used a bootstrap method. Concretely, as described in Section IV-B1, we generated $25$ candidate policies. Subsequently, we sampled five policies with replacements from the $25$ candidates. Through the sampling procedure described above, we randomly made $N=1000$ sets of the five policies.
@@ V Results
V-A Multi-locomotion Domains
The proposed method was evaluated on two Multi-locomotion domains, the Half-Cheetah-Dir domain and the Half-Cheetah-Vel domain, and compared with previous Meta-RL methods, MAML[8], RL2 w/ PPO[34, 2], ProMP [35], PEARL[2]. Figure 5 shows the average returns. As for our proposed method, the average return was derived over five learning trials. The training steps correspond to the foundation policy acquisition phase. For example, since we trained five policies to select one foundation policy, the learning performance at the $500$k time step corresponds to the performance of the foundation policy selected from five candidate policies acquired over $100$k training time steps. For the baseline methods, learning performances were extracted from the data provided in [36]. In the Half-Cheetah-Dir domain, as shown in Fig. 5, our proposed method significantly outperformed all the baseline methods. In the Half-Cheetah-Vel domain, as plotted in Fig. 5, our proposed method showed much better learning performance at the early learning stage, and eventually achieved comparable performance with the PEARL method [2] and outperformed other three baseline methods. Both our proposed and PEARL methods seemed to achieve close to the optimal control performance of the given domain, and the performance of both methods was saturated after about $1$M time steps.
(a) Half-Cheetah-Dir
(b) Half-Cheetah-Vel
Figure 5: Average Returns vs. Training Steps. Error bars show the standard deviation. (a) Proposed method significantly outperformed all the baseline methods. (b) Proposed method showed better learning performance at the early learning stage, and eventually achieved comparable performance with the PEARL method [2] and outperformed other three baseline methods. Both our proposed and PEARL methods seemed to achieve close to the optimal control performance of the given domain, and the performance of both methods was saturated after about $1$M time steps. Proposed method showed comparable performance with the PEARL and outperformed other three baseline methods.
V-B Skill Generation Performance
Domain randomization (DR) is a frequently used approach to cope with multiple tasks and environments. We first evaluated the acquired monopod heading policies using DR as the baseline. We trained the networks under the same combination of context variables as that described in Section IV-B1. Notably, the proposed method also trained the encoder and policy networks, while the DR method did not.
Figures. 6 and 6 show the monopod heading task performances at a novel goal position. As shown in the rightmost panel of Fig. 6, the ball deviated from the goal position when a policy acquired by DR was adopted. Using the proposed method, the ball hit by the monopod successfully tracked the desired trajectory and passed around the goal position, as depicted in Fig. 6. Notably, the monopod model also learned to maintain balance when hitting the ball.
Figure. 6 shows the task performances of the policies trained by the DR (green bar) and the proposed adaptation (red bar). Each box plot was depicted based on $32$ task performances, as described in Section IV-B3. Both the DR and generation phase of the proposed method trained the networks for $3\times 10^{6}$ time steps, which approximately corresponds to $45,000$ trials. The skill-generation phase only required additional $k_{\mathrm{max}}=100$ trials to perform novel tasks successfully.
@@ V-B Skill Generation Performance
Figure 6: Monopod heading performances with the novel goal position. Leftmost snapshot shows the timing of hitting the ball, that is, $t=1.0$ s. Green sphere indicates the goal position. Red dot indicates the desired trajectory. (a) Heading performance by domain randomization. As shown in the rightmost panel, at $t=2.0$ s, the ball deviated from the goal position. (b) Heading performance is performed using the proposed method. Hit ball successfully tracked the desired trajectory and passed around the goal position. Notably, the monopod also learned to maintain its balance with the hit ball. (c) Skill generation performance: domain randomization approach could not achieve a satisfactory multitask performance (green bar). Multiple tasks in multiple environments are accomplished using the proposed method (red bar).
V-C Policy Selection Performance
Here, we evaluate how the selected policy based on the proposed method performed better than a randomly selected model. We adopted the top-one regret as the performance measure to compare the selection performances. Section IV-B4 describes a concrete calculation of the top-one regret.
The top-one regret score of the proposed method was $1.4$, whereas that of the random selection was $5.1$. As the smaller regret showed better performance, we confirmed that the proposed method contributed to selecting a potentially well-performing policy after the adaptation phase. In our approach, we used the performance index described in Eq. (15) for the policy selection. We selected the best performing policy based on the index. Therefore, we did not need to calibrate any threshold. Furthermore, we used the same hyperparameters of the performance index, $\alpha$, $\beta$, and $\gamma$, for all experimental setups. This suggests that our selection approach was robust to changes in the target tasks.
V-D Novel Skill Generation Performance
@@
We depicted how the proposed skill generation method improved the task performance in Fig. 7. As shown in Fig. 7, the hit ball did not track the desired trajectory and deviated from the goal before the adaptation. To evaluate task performance before the skill generation, we set the latent variable $z=0$, assuming that we have prior knowledge of the novel task. After the skill generation, the optimized skill successfully hit the ball towards the new goal position, as shown in Fig. 7.
We also evaluated the overall performances of the policies before and after the skill generation on $32$ different novel tasks, wherein the tasks were sampled randomly, as described in Section IV-B3. Figure 7 shows the task performances of each policy. Before the skill generation, for most novel tasks, task performances were unsatisfactory. Furthermore, performances were varied. After the skill generation, the generated policy for novel tasks consistently achieved successful heading performances.
Furthermore, when we placed the goal position at the back of the robot, the robot model generated a whole new motor skill to “kick” the ball backward to the goal position instead of heading it. Figure 9 illustrated the generated overhead kick movement to hit the ball to the totally new goal position.
The above results indicate that the proposed low-dimensional skill representation and derivative-free optimization successfully facilitate skill emergence in novel situations.
@@
Figure 7: Multitask Heading performance on novel tasks. Green sphere indicates the goal position. Red dot indicates the desired trajectory. (a) Foundational policy: the hit ball did not track the desired trajectory and deviated from the goal. (b) Generated skill: the adapted policy successfully hit the ball toward the new goal position. (c) Multitask heading performances: With the foundational policy, task performances were unsatisfactory and varied. With the generated skill, generated skills for novel tasks consistently achieved successful heading performances.
@@
(b) Skill generation performance
Figure 8: (a) Fine-tuning performance. We found that task performances became worse than those before fine-tuning trials. This result indicated that the policy model could not adapt to new domains with only a small number of additional trials. (b) Proposed method was compared with the baseline method using context variable information. Interestingly, the proposed skill-generation method showed a better heading performance than the method using context information. This result empirically showed that we extracted a meaningful latent space for the given movement category.
Figure 9: Generated overhead kicking skill. We set the goal position (green) in the rear region of the robot, while all training in the foundational policy acquisition phase was done with the goal position in the front region. Thus, here we tried to evaluate the extrapolation performance of our skill generation method. As a result, the robot model generated a whole new motor skill to “kick” the ball backward to the goal position instead of heading it.
@@ VI-A Skill Generation through Latent Variable Optimization
Figure 8 shows the heading performances with the $32$ randomly selected domains. We found that the task performances became worse than those before the fine-tuning trials. This result indicated that the domain randomization policy model could not adapt to new domains with only a small number of additional trials. In other words, extracting the low-dimensional latent space and optimizing the latent variable were significant procedures for generating a skill to cope with novel tasks and environments.
@@ VI-B Using Explicit Context Information
In this study, we assumed that the policy could not observe the context variable in the adaptation phase as we usually lack information regarding novel domains. However, it would be interesting to investigate how task performance can be improved if the context variable was observable. Here, we explicitly encoded the context variable of $32$ novel domains using the trained encoder in the skill-generation phase. The encoder forwards the latent variable to the policy network.
Figure 8 shows the task performances using explicit context information. Interestingly, the proposed adaptation method showed better heading performances than the method using context information. Since the explicit encoder has not been trained in the context of novel domains, the performance degrades. However, the generation skill can adapt to the novel context using the reward signal. Consequently, our proposed method achieved better performance than the method that considers the context variables as inputs. This result empirically showed that we extracted a meaningful latent space for the given movement category.
VI-C Limitations
Our three-phase multitask RL method requires the foundational policy acquisition phase in which multiple candidate policies must be trained. However, the number of required learning trials grows only linearly with the number of candidate policies. Therefore, we have shown that our proposed method exhibits better learning performance than the baseline methods even when explicitly considering the number of learning trials required to train the multiple candidate policies. On the other hand, the number of required candidate policies is still a hyperparameter for our learning system. Although we can parallelize the computation of candidate policy learning, as a future study, we are interested in developing a method to regularize the learning process to guide an acquired policy to be adaptable to generate skills for novel tasks, rather than generating multiple candidate policies and selecting one among the candidates, to further reduce the computational complexity of the learning framework. In this study, we have shown that acquisition of the foundation policy is the key to adaptation to new task goals and environments. However, adaptation to completely different task domains has not been solved even for the large-scale model trained with large-scale data [37, 38]. As a future study, we will investigate whether our proposed approach can deal with a wider range of domains by considering a wider range of contexts in the foundation policy learning with the use of parallel computation in multiple candidate policy learning processes.
@@ VII Conclusion
In motor skill learning, the acquisition of policies that allow rapid adaptation to new situations with different environments and tasks is an important issue. Through the standard multi-locomotion benchmarks and the heading task using the monopod robot, it was verified that the proposed three-phase learning method can quickly generate policies that can cope with multiple tasks. Our framework for learning a foundational policy model based on encoder learning, which embeds explicit contexts in latent spaces, and for selecting and generating motor skills, can efficiently embed and extract the ability to handle multiple tasks. Our proposed method outperformed baseline multitask RL methods on the multi-locomotion benchmarks. Furthermore, by exploiting this skill generation ability in the monopod heading task, we have successfully generated an overhead kicking skill that was not included in the foundational policy acquisition phase at all.
As a future study, we aim to explore the possibility of the proposed method on a multi-agent learning problem. We are interested in observing how our proposed method can modify agent policy according to implicit changes in other agent policies.